• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Carlos

    Log entry time 23:56:21 on October10,2004

    Entry number 133707

    keyword=Swing shift summary

    Shift crew:
    C. Munoz Camacho
    R. Lindgren
    E. Kuchina

    Swing shift summary:

    Very high trip rate at the beginning of the shift. Running smoothly all evening since around 17h.

    16:00 Calorimeter LED scan underway.

    16:15 Asked MCC for beam back (3uA). Resuming DVCS production on LH2, low Q2 kinematics.

    19:00 TWOARMS on LH2 run.

    19:05 Moved target to multifoil.

    19:10 Optics run

    19:15 Back to LH2 target.

    19:20 Restarting DVCS production (3uA).

    19:42 DVCS IOC reboot (see halog entry).

    19:50 Compton computer rebooted (see halog entry).

    00:00 Keeping on with DVCS production.